OpenStudy (anonymous):

A farmer died and left his 11 horses to his 3 sons, such that the eldest was to receive 1/2 of the horses, middle son to receive 1/4 and the youngest to receive 1/6 of the horses. How can they be divided ?

OpenStudy (goformit100):

@QuestionsWolf

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

I don't know how to do this.

OpenStudy (anonymous):

Sorry.

OpenStudy (dumbcow):

im guessing by rounding, since we dont plan on killing horses :) 11/2 = 5.5 --> 6 11/4 = 2.75 --> 3 11/6 = 1.83 --> 2

OpenStudy (jdoe0001):

hehe

OpenStudy (jdoe0001):

1/2 + 1/4 + 1/6 gives 11/12, not 1 so the fractions will give a rational and thus it has to be rounded off

OpenStudy (asnaseer):

add up the fractional amounts to get:\[\frac{1}{2}+\frac{1}{4}+\frac{1}{6}=\frac{6+3+2}{12}=\frac{11}{12}\]so if I lend one horse of my own to make 12 horses altogether, then each one can take 6, 3 and 2 horses respectively, leaving me to take the horse I lent them back. :)

OpenStudy (asnaseer):

this is from a famous problem involving camels

OpenStudy (jdoe0001):

heheh

OpenStudy (anonymous):

to divide the horses bring one of your own, the total now becomes 12, give 6 to the eldest son, 3 to the middle and 2 to the youngest ( 6+ 3 + 2 = 11) and then you ride your horse into the sunset :)